Circular No. 7195

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 E. B. Waltman, Naval Research Laboratory; and F. D. Ghigo, National

Radio Astronomy Observatory, report: “The Green Bank Interferometer (GBI) has observed a flare in the microquasar GRS 1915+105 on June 8. Flux densities peaked near 0.5 Jy at 2 GHz on June 8.3 UT and were declining from 0.25 Jy at 8 GHz. The data are in the public domain and may be accessed via the GBI web page at http://www.gb.nrao.edu.”

SUPERNOVA 1999cm IN UGC 9766 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999cm (cf. IAUC 7190) was obtained by P. Berlind on June 8.2 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) and reveals it to be a type-Ia supernova at maximum light. Cross-correlation of the host galaxy spectrum with absorption and emission-line templates yields a recession velocity of 13 570 km/s for UGC 9766. The photospheric expansion velocity derived from Si II (rest 635.5 nm) is 10 100 km/s. The age of the supernova derived from the spectral features is 1 ± 3 days after B maximum.

COMET C/1999 H3 (LINEAR) Ephemeris extension to IAUC 7152 (elements MPEC 1999-L11):

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 11 17^{h}07.^{m}61 +57^{o}54.^{′}7 3.252 3.556 99.^{o}0 16.^{o}4 13.1
21 16 29.13 +57 56.8 3.277 3.541 96.6 16.6 13.1
July 1 15 53.83 +56 51.8 3.328 3.528 92.8 16.7 13.1
11 15 24.42 +54 55.8 3.401 3.518 88.1 16.8 13.1
21 15 01.68 +52 27.9 3.491 3.510 82.7 16.7 13.2
31 14 45.03 +49 44.8 3.591 3.504 77.0 16.4 13.2
Aug. 10 14 33.39 +46 58.7 3.696 3.501 71.1 15.9 13.3
20 14 25.71 +44 17.5 3.801 3.500 65.2 15.2 13.3
30 14 21.05 +41 46.1 3.901 3.502 59.7 14.4 13.4

Visual m_{1} estimates: May 22.03 UT, 13.4 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m reflector); 24.89, 12.9 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); June 3.89, 13.6 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 4.98, 12.2 (J. Carvajal, Madrid, Spain, 0.32-m reflector); 7.94, 13.0 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector).

1999 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green